OpenStudy (anonymous):

Solve the system, using substitution. One point for correct steps, one point for correct ordered pair, (a,b). a + b = 4 and 2a + 3b = 11

hero (hero):

a + b = 4 ---> a = 4 - b 2a + 3b = 11 ----> a = (11 - 3b)/2 a = a 2(4-b) = 11 - 3b 8 - 2b = 11 - 3b 3b - 2b = 11 - 8 b = 3 a = 1 (a,b) = (1,3)